write in molecular and ionic form the reaction equation proving the amphoteric properties of the following hydroxides : a. Cr(OH)3 , b. Al(OH)3. what analytical group do these cations belong to ?
1
Expert's answer
2020-01-17T07:01:32-0500

Amphoteric hydroxides are chemicals that can exhibit both basic and acidic properties (depending on reaction conditions). Amphoteric hydroxides behave as bases in an acidic environment, and as acids in an alkaline environment. Amphoteric hydroxides include bases formed by transition metals: Al (3+), Be (2+), Fe (3+), Cr (3+), Zn (2+), Sn (2+), Pb (2+ ) other. Almost all of them are insoluble in water, are weak electrolytes.

When heated, the compounds decompose.

a) Cr(OH)3 + 3HCl -> CrCl3 + 3H2O

Cr(OH)3+3H(+)+3Cl(-)=Cr(3+)+3CL(-)+3H2O

Cr(OH)3+3H(+)=Cr(3+)+3H2O


Cr(OH)3 + 3NaOH -> Na3[Cr(OH)6]

Cr(OH)3+ 3Na(+ )+3OH(-)= 3Na(+)+{ Cr(OH)6}(3-)

Cr(OH)3+3OH(-)={ Cr(OH)6}(3-)


b) Al(OH)3 + 3HCl-> AlCl3 +3H2O

Al(OH)3+3H(+)+3Cl(-)=Al3(+)+3CL(-)+3H2O

Al(OH)3+3H(+)=Al(3+)+3H2O


Al(OH)3+ 3NaOH-> Na3AlO3+ 3H2O

Al(OH)3+ 3Na(+) +3OH(-)= 3Na(+)+ AlO3(3-)+ 3H2O

Al(OH)3+3OH(-)=AlO3(3-)+ 3H2O


These are cations 4 of the analytical group.